OpenStudy (anonymous):

The product of two consecutive positive even integers is 12,320. What are the two integers

OpenStudy (anonymous):

guess and check for this one

Parth (parthkohli):

Why? Why not the great old algebra?

Parth (parthkohli):

\(\Rightarrow x(x + 2) = 12,320 \) \(\Rightarrow x^2 + 2x - 12,320 = 0\)

Parth (parthkohli):

Seems like you need Wolfram..

OpenStudy (anonymous):

ok here is why not the great old algebra

OpenStudy (anonymous):

solving \(x^2+2x-12320=0\) by factoring is identical to solving the original problem, guess at two numbers that are two apart,whose product is 12320

Parth (parthkohli):

I am failing at Mathematics so Idk. 8th grade math is hard D:

OpenStudy (anonymous):

lets try 124 and 126 \(124\times 126=15,624\) ok that is not it, try a smaller pair

Parth (parthkohli):

Wow, my guess worked at the first time. I am not telling what I got as it would be great for fiends.

OpenStudy (anonymous):

i guess one hint is that they are both even and it ends in a zero so a good chance that one of the two numbers ends in a zero as well

OpenStudy (anonymous):

\(118\times 120=14160\) nope still too big try \(110\)

Parth (parthkohli):

The superior Satellite ^

ganeshie8 (ganeshie8):

one more hint : since the product somewhat large, in @ParthKohli equation, \(x^2\) term dominates. a good first guess would be around square root of the product

OpenStudy (anonymous):

\[2n(2n+2)=12320\]\[4 n^2+4 n-12320=0 \]\[4 (n-55) (n+56)=0 \]n=55\[\{2*55,(2*55+2)\} \text{ -$>$} \{110,112\}, 110*112 = 12320 \]